OpenStudy (anonymous):

if you start a job that pays $0.05 the first day, $0.10 the second day and $0.20 the third day and so on, what will your total earnings be after Two weeks?

OpenStudy (anonymous):

0.05+0.10+0.20+0.40....etc etc=919.15 if my math was right

OpenStudy (anonymous):

thats not the correct answer unfortunatley :/ it's $819.15 i can't figure out the difference ratio since they give me two differences (0.05 and 0.10). is there any other way we can figure this out?

OpenStudy (anonymous):

it looks like it just doubles the previous number every time

OpenStudy (anonymous):

so my d=2?

OpenStudy (anonymous):

is 2d a option?

OpenStudy (jdoe0001):

a common ratio of 2 is about right

OpenStudy (anonymous):

no, i'm trying to plug the things i know into a formula the formula is : an=a1=(n-1)d a1=0.05 and d is the difference ratio. i can create my own formula and plug in 14 for n.. i hope that makes sense. @jdoe0001 i'm getting a negative an! which is bad since we're talking paying someone here. suggestions?

OpenStudy (jdoe0001):

hmm

OpenStudy (anonymous):

my formula for an looks like: 2n-1.95 does that look right? @jdoe0001 ?

OpenStudy (jdoe0001):

\(\large \begin{array}{llll} \small1&\small2 &\small 3&...\\ 5 & 10 & 20&...\\ 5^{n-1}&5^{n-1} & ... \end{array}\)

OpenStudy (jdoe0001):

so the 14th member in the geometric sequence will be \(\bf \large a_{14} = a_1\times r^{n-1}\)

OpenStudy (texaschic101):

this is a geometric sequence... sn = a1(1 - r^n)/ (1 - r) a1 = .05 r = 2 lets sub s14 = .05(1 - 2^14)/ (1 - 2) s14 = .05 (1 - 16384) / -1 s14 = .05 (1 - 16384)/-1 s14 = .05(-16383)/-1 s14 = -819.25/-1 s14 = 819.25

OpenStudy (anonymous):

ohhh! i've been using the wrong s formula!!!! thank you so much everyone! @texaschic101 :)

OpenStudy (jdoe0001):

but you only need the sum.. not sure if you need the 14th member for the sum

OpenStudy (texaschic101):

no problem....when you have to multiply to find the difference, then it is a geometric sequence, and when you have to add to find the difference, then it is an arithmetic sequence. And they have different formulas.

OpenStudy (jdoe0001):

using the common ratio of 5 => \(\bf s_{14} = a_1\left(\cfrac{1-5^{14}}{1-5}\right)\) as shown above by texaschic101

OpenStudy (jdoe0001):

anyhohw, I used 5 since I was using cents as based, no $ by the way 0.05 is a rip off :P

OpenStudy (anonymous):

Thanks guys! @jdoe0001 and @texaschic101 :)